OpenStudy (anonymous):

if 4/7 of a circle is shaded and another 1/3 of it is also shaded..how much of it is not shaded?

OpenStudy (anonymous):

Make the fractions have a common denominator and add: 4/7 = 12/21 1/3 = 7/21 Total shaded = 19/21 Not shaded = 1 - (19/21) = 2/21

OpenStudy (anonymous):

yes! thank u!! :)